OpenStudy (anonymous):

what is the length(magnitude) of the vector (6,-3)? A. 3sqrt2 B. sqrt2 C. 3 D. sqrt53

OpenStudy (zzr0ck3r):

length of vector \((a,b)\) is \(\sqrt{a^2+b^2}\)

OpenStudy (mathstudent55):

You need the hypotenuse of the right triangle. 

OpenStudy (anonymous):

would I be using the Pythagorean theorem?

OpenStudy (rhr12):

√(6^2 + (-3)^2) = √(36+9) = √(9*5) = 3√5

OpenStudy (sepeario):

Yes, Pythagoras' Theorem is the way to figure out the length of this vector.